# Semi-Supervised Learning for Lensed Quasar Detection

### v1

@@ Abstract
Lensed quasars are key to many areas of study in astronomy, offering a unique probe into the intermediate and far universe. However, finding lensed quasars has proved difficult despite significant efforts from large collaborations. These challenges have limited catalogues of confirmed lensed quasars to the hundreds, despite theoretical predictions that they should be many times more numerous. We train machine learning classifiers to discover lensed quasar candidates. By using semi-supervised learning techniques we leverage the large number of potential candidates as unlabelled training data alongside the small number of known objects, greatly improving model performance. We present our two most successful models: (1) a variational autoencoder trained on millions of quasars to reduce the dimensionality of images for input to a gradient boosting classifier that can make accurate predictions and (2) a convolutional neural network trained on a mix of labelled and unlabelled data via virtual adversarial training. These models are both capable of producing high-quality candidates, as evidenced by our discovery of GRALJ140833.73+042229.98. The success of our classifier, which uses only images, is particularly exciting as it can be combined with existing classifiers, which use other data than images, to improve the classifications of both models and discover more lensed quasars.
@@
pubyear: 2015pagerange: Semi-Supervised Learning for Lensed Quasar DetectionA
@@ 1 Introduction
Figure 1: Six images of lensed quasars showing data from DESI. The top two images show the archetypal lensed quasar as it would be described to a student: two identically blue-white quasars on either side of a red lensing galaxy, which is often extended. The middle two images show examples of quadruply lensed quasars, four blue-white quasars arranged in a kite. Sometimes, as in the left image, the lensing galaxy is barely or not at all visible and other times, as in the right image, the lensing galaxy obscures one of the quasar images. The bottom two images show more typical lensed quasars, two blue-white quasars, often with the lensing galaxy too dim to be seen and commonly one of the quasar images is reddened, presumably by the lensing galaxy.
Since their first discovery by Walsh et al. (1979), lensed quasars have been hotly sought after astronomical objects as probes of cosmology and the structure of galaxies and their surrounds. Because the light from each image of the lensed quasar has travelled different paths through space-time, although originating at a single source, we can compare the observed spectra of the different images to infer properties of the circumgalactic medium of the lensing galaxy (e.g. Rauch et al., 2001; Rauch & Haehnelt, 2011), the intergalactic medium (e.g. Foltz et al., 1984; Smette et al., 1995) and even the Hubble constant (e.g. Refsdal, 1964; Meyer et al., 2023).
@@
To reliably identify quasars several challenges must be overcome. Firstly, quasars themselves are rare, appearing in only a small fraction of galaxies; lensed quasars are many times rarer, occurring at a rate of about 1/1000110001/10001 / 10001/10 0001100001/10\,0001 / 10 000 quasars. This results in a massive class imbalance for any manual or automated classification effort.
The second difficulty is the small number of confirmed lensed quasars, now sitting at about 250 (e.g. Ducourant et al., 2018; Lemon et al., 2019; Krone-Martins et al., 2019), with perhaps 400 more which have been identified but as yet are unpublished. Compared to modern image recognition tasks, which often utilise billions of images, this number is painfully small; even the number of known unlensed quasars only sits at around 6 million (Flesch, 2023; Gaia Collaboration et al., 2023). Labelling more data is also prohibitively expensive: to confirm whether an object is a lensed quasar we require a 3.5+ m telescope operated for hours by a highly skilled astronomer, which can still lead to an inconclusive result. Thus, while a handful of such observations can be performed each year, in the short term our set of lensed quasars will remain of size $\sim$650.
@@
Another difficulty is the quality of the available data: not all of the objects listed in catalogues like Milliquas (Flesch, 2023) are quasars. “Quasars” in catalogues like Milliquas have often been identified by applying machine learning to large surveys and, although often the best course of action, leads to mislabelled objects. Images of these quasars can be taken from the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; Chambers et al., 2016) survey (covering the northern sky) and/or the Dark Energy Spectroscopic Instrument (DESI; Dey et al., 2019) Legacy Imaging Surveys, covering the southern sky. Image quality and appearance differ between these two surveys and visual artefacts can further complicate classification of lensed quasars. Compared to image classification tasks in other domains, the images of quasars we work with are often significantly more noisy, as shown in Figure 2. To further complicate matters, lensed quasars are partially absent from surveys such as Gaia. Despite Gaia’s exceptional resolving power of 180 mas, some quasar images are absent from current data releases because they are too close together, resulting in poorer astrometry which does not pass the filters selecting the sources entering the catalogue. Thus, there is likely a fraction of lensed quasars which will remain unidentified if naively performing analysis on surveys like Gaia (even after exploiting subtle signatures, e.g. Mannucci et al., 2022).
@@
The final problem is that undiscovered lensed quasars likely have different properties to the discovered population of lensed quasars, or else they would have been discovered in previous searches. Concretely, undiscovered lensed quasars likely have a smaller separation, have an image which is reddened or have an image disguised in some other fashion. This adds another challenge to the classification problem: ideally we need a model which can generalise from our small training set to select candidates which are from a different distribution than discovered lensed quasars. This need to generalise breaks independent and identically distributed assumptions (that a typical candidate is drawn from the same distribution as the training data) which underpin most machine learning techniques.
@@ 2.1 Lensed Quasars
Our list of known lensed quasars comes from a catalogue by Krone-Martins (priv. comm., 2021) which includes some unpublished lensed quasars. In this catalogue the coordinates for each image are listed separately. To avoid cluttering the data set with very similar images, one image from each lensed quasar was randomly selected to be included and the rest were discarded.
@@ 2.3 Pan-STARRS
In this work we draw images from two surveys, the first of which is the Pan-STARRS1 survey DR2 (Magnier et al., 2020; Waters et al., 2020; Flewelling et al., 2020). The Pan-STARRS survey faces the northern sky, imaging north of $-30^{\circ}$ of declination in five broadband filters: g, r, i, z, y. Pan-STARRS images are often noisy, as shown in Figure 2, but in general do provide g, r and i bands. The images can be downloaded from the University of Hawaii via their Image Cutout Service.
Unfortunately, downloading data from the servers as FITS files is quite slow with the download time increasing linearly with the number of bands requested. However, in the time taken to download a single band, a 3-colour PNG or JPEG image can be downloaded, allowing us to effectively triple the download rate. We experimented with PNG and JPEG formats and found that the JPEG compression enhanced model performance by smoothing the high-frequency noise in the images. We downloaded 3.8 million 64x64 pixel JPEG images, corresponding to 16x16 arcseconds of sky, in the g, r and i bands around the quasars and lensed quasars described in Sections 2.1 and 2.2.
@@ 2.4 DESI Legacy Surveys
Due to its location in Hawaii, USA, the Pan-STARRS survey primarily images the northern sky, so we require another survey to image quasars in the southern sky. The DESI Legacy Imaging Survey DR10 (henceforth DESI survey) has good coverage south of $-20^{\circ}$ of declination, but often only in select bands. To match the Pan-STARRS images we again chose 64x64 JPEG images in the g, r and i bands, this time numbering 2 million fields. Unfortunately, the DESI survey has not achieved full coverage in the g, r and i bands, so some of our DESI images are 2-colour. In order to handle these null values, we opted to fill the values with zeros and pass a flag to the classifiers to indicate in which colour bands null values existed in the original image. Unlike the Pan-STARRS data, DESI survey data can be quickly downloaded for several channels, allowing us to download FITS files for each band and create the JPEGs locally. We found this resulted in less blurred JPEG images compared to JPEGs downloaded directly from the DESI survey servers.
Figure 3: Diagram showing the architecture for the autoencoder-classifier model where z depicts that latent space of the autoencoder. The autoencoder architecture and training is described in Sections 3.2 and 3.3. The noise metric is described in Section 3.4.1. The traditional classifiers and metadata are described in Section 3.4.2.
@@ 3 Methods
Machine learning is finding ever increasing use in astronomy. In the modern era of huge numbers of real time science alerts, machine learning is one of a few tools capable of triaging the deluge of astronomical data collected by our telescopes. Particularly relevant to this work are the studies performed by data brokers such as ALeRCE (Förster et al., 2021) to classify low resolution data. Typically these models are trained on moderate sets of labelled data to classify the stream of data and dispatch science alerts to interested astronomers (Carrasco-Davis et al., 2021; Sánchez-Sáez et al., 2021a) or to identify anomalies in the data (Sánchez-Sáez et al., 2021b; Perez-Carrasco et al., 2023). They have developed a range of novel techniques, for example, segmenting an image and providing it to a model at different scales (Förster et al., 2022). These studies provided us with inspiration for our model architectures and we hope the present work will in turn spark further new avenues in future.
@@ 3.2 Convolutional Autoencoder
Figure 5: Reconstructions of two (previously unseen by the model) lensed quasars generated by autoencoders with varying bottlenecks, including the best performing autoencoder (described at the end of Section 3.4). Note that as the dimensionality of the latent space increases the reconstruction fidelity improves, however this often does not translate to increased performance for the classifier as the increased dimensionality of the latent space can reduce performance.
Autoencoders (Hinton & Salakhutdinov, 2006) are deep neural networks with an hourglass structure which are trained to recreate the input, in our case an image, with their output layer. The initial layer, with the same dimensions as the image, is followed by successively smaller layers until a “bottleneck” layer is reached — this part of the network is the encoder. After the bottleneck layer the structure is reversed, with each successive layer getting larger until the original dimensionality is reached — called the decoder. The bottleneck layer typically has small dimensionality (of order 10) which forces the network to reduce the information content of the image into a few values before reconstituting the image. Reducing the dimensionality of the bottleneck layer allows less information to be passed to the decoder, resulting in a coarser reconstruction (as shown in Figure 5), but a more information dense latent space. By training a network in this way we obtain an encoder which is capable of reducing the dimensionality of the input image so that the classification task is tractable for a traditional classifier (e.g. a random forest; Breiman, 2001).
@@ 3.3 $\beta$-Variational Autoencoder
We also trained $\beta$-variational autoencoder models (Higgins et al., 2017) which performed better, i.e. resulted in superior classifications compared to the simple autoencoder models. Variational autoencoders (VAEs; Kingma & Welling, 2013) have the same underlying structure as autoencoder models but with the addition of two major elements. The first is that instead of connecting the encoder and decoder directly, the autoencoder outputs a mean and standard deviation for each of the “usual” bottleneck neurons. The decoder then samples from a normal distribution using the means and standard deviations provided and proceeds with these sampled values. The second difference is that the loss function, given in Equation 1, has an added term which penalises the network based on the Kullback-Leibler (KL) divergence (Kullback & Leibler, 1951) between the normal distributions described by the means and standard deviations and normal distributions with 0 means and standard deviations of 1. These modifications to the autoencoder encourages it to map images into a latent space at the bottleneck layer which is more meaningful. The random sampling at the bottleneck layer makes it much more difficult for the autoencoder to “memorise” training examples by mapping them to specific locations in the latent space, while the added penalty to the loss function ensures the means and standard deviations themselves are hindered from overfitting. Thus, the loss function, $\mathcal{L}(y,\hat{y})$, is:
$\mathcal{L}(y,\hat{y})$ $=\text{MSE}(y,\hat{y})+\beta\cdot D_{\text{KL}}\left(\mathcal{N}(0,1)\,||\,\mathcal{N}(\mu,\sigma^{2})\right)$ (1)
Where $\text{MSE}(y,\hat{y})$ is the mean squared error between the images and their reconstructions by the model and $D_{\text{KL}}\left(\mathcal{N}(0,1)\,||\,\mathcal{N}(\mu,\sigma^{2})\right)$ is the KL divergence between the means and standard deviations produced by the autoencoder and the standard normal distribution.
@@ 3.4.2 Traditional Classifiers
We trained a number of traditional models to classify whether objects were lensed quasars. These classifiers were provided with the latent space values of the chosen autoencoder, the reconstruction errors, the noise metrics, whether each image was from the Pan-STARRS or DESI catalogue and if from the DESI catalogue whether the g, r and i channels were null. For every autoencoder model, 5-fold cross validation tuning was performed to identify the best set of hyperparameters for each classifier. This resulted in classifiers trained for autoencoder and $\beta$-VAE models with bottlenecks of 4/8/16/32/64 and, in the case of the $\beta$-VAE models, various values of $\beta$. The classifiers we chose to train were: random forest (Breiman, 2001), isolation forest (Liu et al., 2012), logistic regression (Berkson, 1944), gradient boosting (Friedman, 2001), support vector machine (SVM; Vapnik, 1995), kernel-SVM (Cortes & Vapnik, 1995) and Gaussian process (Rasmussen & Williams, 2006).
The 5-fold cross validation tuning comprised randomly searching a predefined space of hyperparameters using scikit-learn’s RandomizedSearchCV and selecting the set of hyperparameters which achieved the highest F1 score. The details of the cross validation tuning of each classifier are in Appendix A. All models and the random search were seeded with a value of 0, so the results may be recreated.
The best autoencoder-classifer combination achieved an F1 score of 0.87 by using a $\beta$-VAE with $\beta=0.0001$ (larger and smaller $\beta$ values were tested) and a bottleneck of 8 combined with a gradient boosting classifier having a subsample of 1, a ‘scale_pos_weight’ of 105.37, 249 estimators, a maximum depth of 21 and a learning rate of 1.889. The performance of tuned classifiers combined with $\beta$-VAEs with $\beta=0.0001$ are shown in Figure 7.
Figure 7: Performance of the traditional classifiers on $\beta$-VAE models with $\beta=0.0001$ for various bottlenecks. The highest F1 score was achieved by an autoencoder-classifier combination employing a $\beta$-VAE model with $\beta=0.0001$ and bottleneck $=8$ and the gradient boosting classifier.

### v2

@@ Abstract
Lensed quasars are key to many areas of study in astronomy, offering a unique probe into the intermediate and far universe. However, finding lensed quasars has proved difficult despite significant efforts from large collaborations. These challenges have limited catalogues of confirmed lensed quasars to the hundreds, despite theoretical predictions that they should be many times more numerous. We train machine learning classifiers to discover lensed quasar candidates. By using semi-supervised learning techniques we leverage the large number of potential candidates as unlabelled training data alongside the small number of known objects, greatly improving model performance. We present our two most successful models: (1) a variational autoencoder trained on millions of quasars to reduce the dimensionality of images for input to a dense neural network classifier that can make accurate predictions and (2) a convolutional neural network trained on a mix of labelled and unlabelled data via virtual adversarial training. These models are both capable of producing high-quality candidates, as evidenced by our discovery of GRALJ140833.73+042229.98. The success of our classifier, which uses only multi-band images, is particularly exciting as it can be combined with existing classifiers, which use other data than images, to improve the classifications of both models and discover more lensed quasars.
@@
pubyear: 2015pagerange: Semi-Supervised Learning for Lensed Quasar DetectionC
@@ 1 Introduction
Figure 1: Six images of lensed quasars showing data from DESI. The top two images show the archetypal lensed quasar as it would be described to a student: two identically blue-white quasars on either side of a red lensing galaxy, which is often extended. This typical difference in colouring between the quasar and lensing galaxy is the reason multi-band images are so crucial — to differentiate between other astrophysical phenomena. The middle two images show examples of quadruply lensed quasars, four blue-white quasars arranged in a kite. Sometimes, as in the left image, the lensing galaxy is barely or not at all visible and other times, as in the right image, the lensing galaxy obscures one of the quasar images. The bottom two images show more typical lensed quasars, two blue-white quasars, often with the lensing galaxy too dim to be seen and commonly one of the quasar images is reddened, presumably by the lensing galaxy. Note that the images in this paper are comprised of the $g$, $r$ and $i$ bands and visualised in images as the r, g, b channels of a standard image. Unless otherwise specified, the images are 16x16 arcsec of sky.
Since their first discovery by Walsh et al. (1979), lensed quasars have been hotly sought after astronomical objects as probes of cosmology and the structure of galaxies and their surrounds. Because the light from each image of the lensed quasar has travelled different paths through space-time, although originating at a single source, we can compare the observed spectra of the different images to infer properties of the circumgalactic medium of the lensing galaxy (e.g. Rauch et al., 2001; Rauch and Haehnelt, 2011), the intergalactic medium (e.g. Foltz et al., 1984; Smette et al., 1995) and even the Hubble constant (e.g. Refsdal, 1964; Meyer et al., 2023).
@@
To reliably identify quasars several challenges must be overcome. Firstly, quasars themselves are rare, appearing in only a small fraction of galaxies; lensed quasars are many times rarer, occurring at a rate of about 1/10001/10001/10 0001/10\,000 quasars. This results in a class imbalance for any manual or automated classification effort.
The second difficulty is the small number of confirmed lensed quasars, now sitting at about 250 (e.g. Ducourant et al., 2018; Lemon et al., 2019; Krone-Martins et al., 2019), with perhaps 400 more which have been identified but as yet are unpublished. Compared to modern image recognition tasks, which often utilise billions of images, this number is painfully small; even the number of known quasars (which may or may not be lensed) only sits at around 6 million (Flesch, 2023; Gaia Collaboration et al., 2023). Labelling more data is also prohibitively expensive: to confirm whether an object is a lensed quasar we require a 3.5+ m telescope operated for hours by a highly skilled astronomer, which can still lead to an inconclusive result. Thus, while a handful of such observations can be performed each year, in the short term our set of lensed quasars will remain of size $\sim$650.
@@
Another difficulty is the quality of the available data: not all of the objects listed in catalogues like Milliquas (Flesch, 2023) are quasars. “Quasars” in catalogues like Milliquas have often been identified by applying machine learning to large surveys and, although often the best course of action, leads to mislabelled objects. Images of these quasars can be taken from the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; Chambers et al., 2016) survey (covering the northern sky) and/or the Dark Energy Spectroscopic Instrument (DESI; Dey et al., 2019) Legacy Imaging Surveys, covering the southern sky. Image quality and appearance differ between these two surveys and visual artefacts can further complicate classification of lensed quasars. Compared to image classification tasks in other domains, the images of quasars we work with are often significantly more noisy, as shown in Figure 2. To further complicate matters, lensed quasars are partially absent from surveys such as Gaia. Despite Gaia’s exceptional resolving power of 180 mas, some quasar images are absent from current data releases because they are too close together, resulting in poorer astrometry which does not pass the filters selecting the sources entering the catalogue. Thus, there is likely a subset of lensed quasars which are absent from surveys like Gaia (even after exploiting subtle signatures, e.g. Mannucci et al., 2022), despite having been observed.
@@
The final problem is that undiscovered lensed quasars likely have different properties to the discovered population of lensed quasars, or else they would have been discovered in previous searches. Concretely, undiscovered lensed quasars likely have a smaller separation, have an image which is reddened or have an image disguised in some other fashion. This adds another challenge to the classification problem: ideally we need a model which can generalise from our small training set to select candidates which are from a different distribution than discovered lensed quasars. This need to generalise breaks assumptions which underpin most machine learning techniques (e.g. that the data are independent and identically distributed across the training and unseen data).
@@ 2.1 Lensed Quasars
Our list of known lensed quasars comes from a catalogue by Krone-Martins (priv. comm., 2021). This catalogue is a collation of published lensed quasars as well some some as yet unpublished lensed quasars identified by GraL collaboration (Ducourant, submitted for publication). In this catalogue the coordinates for each image are listed separately. To avoid cluttering the data set with very similar images, one image from each lensed quasar was randomly selected to be included and the rest were discarded.
@@ 2.3 Pan-STARRS
In this work we draw images from two surveys, the first of which is the Pan-STARRS1 survey DR2 (Magnier et al., 2020; Waters et al., 2020; Flewelling et al., 2020). The Pan-STARRS survey faces the northern sky, imaging north of $30^{\circ}$ of declination in five broadband filters: g, r, i, z, y. Pan-STARRS images are often noisy, as shown in Figure 2, but in general do provide g, r and i bands. The images can be downloaded from the University of Hawaii via their Image Cutout Service.
Unfortunately, downloading data from the servers as FITS files is quite slow with the download time increasing linearly with the number of bands requested. However, in the time taken to download a single band, a 3-colour PNG or JPEG image can be downloaded, allowing us to effectively triple the download rate (note that recent upgrades to the hosting of the Pan-STARRS data have rendered this workaround unnecessary). We experimented with PNG and JPEG formats and found that the JPEG compression (see Appendix A for details) enhanced model performance by smoothing the high-frequency noise in the images. We downloaded 3.8 million 64x64 pixel JPEG images, corresponding to 16x16 arcseconds of sky, in the g, r and i bands around the quasars and lensed quasars described in Sections 2.1 and 2.2.
@@ 2.4 DESI Legacy Surveys
Due to its location in Hawaii, USA, the Pan-STARRS survey primarily images the northern sky, so we require another survey to image quasars in the southern sky. The DESI Legacy Imaging Survey DR10 (henceforth DESI survey) has good coverage south of $20^{\circ}$ of declination, but often only in select bands. To match the Pan-STARRS images we again chose 64x64 JPEG images in the g, r and i bands, this time numbering 2 million fields. Unfortunately, the DESI survey has not achieved full coverage in the g, r and i bands, so some of our DESI images are 2-colour. In order to handle these null values, we opted to fill the values with zeros and pass a flag to the classifiers to indicate in which colour bands null values existed in the original image. Unlike the Pan-STARRS data, DESI survey data can be quickly downloaded for several channels, allowing us to download FITS files for each band and create the JPEGs locally. We found this resulted in less blurred JPEG images compared to JPEGs downloaded directly from the DESI survey servers.
Figure 3: Diagram showing the architecture for the autoencoder-classifier model where z depicts that latent space of the autoencoder. The autoencoder architecture and training is described in Sections 3.2 and $3.$ The noise metric is described in Section 3.4.1. The traditional classifiers and metadata are described in Section 3.4.2.
@@ 3 Methods
Machine learning is finding ever increasing use in astronomy. In the modern era of huge numbers of real time science alerts, machine learning is one of a few tools capable of triaging the deluge of astronomical data collected by our telescopes. Particularly relevant to this work are the studies performed by data brokers such as ALeRCE (Förster et al., 2021) to classify low resolution data. Typically these models are trained on moderate sets of labelled data to classify the stream of data and dispatch science alerts to interested astronomers (Carrasco-Davis et al., 2021; Sánchez-Sáez et al., 2021b) or to identify anomalies in the data (Sánchez-Sáez et al., 2021a; Perez-Carrasco et al., 2023). They have developed a range of novel techniques, for example, segmenting an image and providing it to a model at different scales (Förster et al., 2022). These studies provided us with inspiration for our model architectures and we hope the present work will in turn spark further new avenues in future.
@@ 3.2 Convolutional Autoencoder
Figure 5: Reconstructions of two (previously unseen by the model) lensed quasars generated by autoencoders with varying bottlenecks, including the best performing autoencoder (described at the end of Section 3.4). Note that as the dimensionality of the latent space increases the reconstruction fidelity improves, however this often does not translate to increased performance for the classifier as the increased dimensionality of the latent space can reduce performance. These images represent 8x8 arcsec of sky.
Autoencoders (Hinton and Salakhutdinov, 2006) are deep neural networks with an hourglass structure which are trained to recreate the input, in our case an image, with their output layer. The initial layer, with the same dimensions as the image, is followed by successively smaller layers until a “bottleneck” layer is reached — this part of the network is the encoder. After the bottleneck layer the structure is reversed, with each successive layer getting larger until the original dimensionality is reached — called the decoder. The bottleneck layer typically has small dimensionality (of order 10) which forces the network to reduce the information content of the image into a few values before reconstituting the image. Reducing the dimensionality of the bottleneck layer allows less information to be passed to the decoder, resulting in a coarser reconstruction (as shown in Figure 5), but a more information dense latent space. By training a network in this way we obtain an encoder which is capable of reducing the dimensionality of the input image so that the classification task is tractable for a traditional classifier (e.g. a random forest; Breiman, 2001).
@@ 3.3 $\beta$-Variational Autoencoder
[…]
@@ 3.4.2 Traditional Classifiers
We trained a number of traditional models to classify whether objects were lensed quasars. These classifiers were provided with the latent space values of the chosen autoencoder, the reconstruction errors, the noise metrics, whether each image was from the Pan-STARRS or DESI catalogue and if from the DESI catalogue whether the g, r and i channels were null. For every autoencoder model, 5-fold cross validation tuning was performed to identify the best set of hyperparameters for each classifier. This resulted in classifiers trained for autoencoder and $\beta$-VAE models with bottlenecks of 4/8/16/32/64 and, in the case of the $\beta$-VAE models, various values of $\beta$. The classifiers we chose to train were: random forest (Breiman, 2001), isolation forest (Liu et al., 2012), logistic regression (Berkson, 1944), gradient boosting (Friedman, 2001), support vector machine (SVM; Vapnik, 1995), kernel-SVM (Cortes and Vapnik, 1995), Gaussian process (Rasmussen and Williams, 2006) and a densely connected artificial neural network.
The 5-fold cross validation tuning for all models except the neural network comprised randomly searching a predefined space of hyperparameters using scikit-learn’s RandomizedSearchCV and selecting the set of hyperparameters which achieved the highest F1 score (defined in Appendix B). The neural network was tuned using scikit-learn’s StratifiedKFold and a grid search over the hyperparameters. The details of the cross validation tuning of each classifier are in Appendix C. All models and the random search were seeded with a value of 0, so the results may be recreated. Finally, models were compared using stratified 5-fold cross validation using StratifiedKFold with the median F1, precision and recall scores shown in Figure 7
The best autoencoder-classifer combination achieved an F1 score of 0.90 by using a $\beta$-VAE with $\beta=0.0001$ (larger and smaller $\beta$ values were tested) and a bottleneck of 32 combined with a densely connected artificial neural network with two hidden layers of 32 nodes and a weight decay of 0.003. The performance of tuned classifiers combined with $\beta$-VAEs with $\beta=0.0001$ are shown in Figure 7.
Figure 7: Performance of the traditional classifiers on $\beta$-VAE models with $\beta=0.0001$ for various bottlenecks. The highest F1 score was achieved by an autoencoder-classifier combination employing a $\beta$-VAE model with $\beta=0.0001$ and bottleneck =8=8 and the gradient boosting classifier.